### v1

# Thermodynamic Limit in Learning Period Three

Abstract
A continuous one-dimensional map with period three includes all periods. This raises the following question: Can we obtain any types of periodic orbits solely by learning three data points? We consider learning period three with random neural networks and report the universal property associated with it. We first show that the trained networks have a thermodynamic limit that depends on the choice of target data and network settings. Our analysis reveals that almost all learned periods are unstable and each network has its characteristic attractors (which can even be untrained ones). Here, we propose the concept of characteristic bifurcation expressing embeddable attractors intrinsic to the network, in which the target data points and the scale of the network weights function as bifurcation parameters. In conclusion, learning period three generates various attractors through characteristic bifurcation due to the stability change in latently existing numerous unstable periods of the system.
With the advent of reservoir computing (RC) [1], which exploits the dynamics of high-dimensional dynamical systems for learning, many interesting properties of random networks have been discovered from a dynamical systems perspective. For example, with a fixed-weight random recurrent neural network, the so-called echo-state network (ESN) [2]—as a reservoir—we can create an autonomous system that emulates target chaotic systems by simply fitting its readout layer [3, 4]. Such an attractor-embedding ability is linked to the existence of the generalized synchronization between an input dynamical system and a reservoir [4, 5, 6, 7]. In addition, recent studies have revealed that RC can simultaneously embed multiple attractors [8] and may have untrained attractors that are not part of the training data [8, 9, 10, 11, 12]. Surprisingly, a successful reservoir computer only needs a few pairs of bifurcation parameter values and corresponding trajectories to reconstruct the entire bifurcation structure of a target system [10, 11]. These properties are valuable, for example, in the context of robot locomotion control using dynamical system attractors that can significantly reduce the training data [13, 14, 15]. Thus, the powerful generalization and multifunctionality aspects of RC are related to the dynamical systems properties of a learning machine.
In one-dimensional discrete dynamical systems, there are two significant theorems on periodic orbits [16, 17, 18]:
Theorem 1
(Sharkovsky) If a continuous map $f:I\rightarrow I$ has a periodic point of period $m$ and $m\succ n$, then $f$ also has a periodic point of period $n$.
Theorem 2
(Li–Yorke) If a continuous map $f:I\rightarrow I$ has a point $a\in I$ for which the points $b=f(a)$, $c=f(f(a))$, and $d=f(f(f(a)))$ satisfy
$d\leq a<b<c\quad({\rm or}\quad d\geq a>b>c),$
then $f$ has a periodic point of period $k$ for every $k\in\mathbb{N}$.
Note that the interval $I$ does not need to be closed or bounded. Here, the ordering of positive integers $\succ$ in Theorem 1 is called the Sharkovsky ordering and is given below:
33 33 $\succ 5\succ 7\succ 9\succ\cdots\succ 2\cdot 3\succ 2\cdot 5\succ 2\cdot 7\succ\cdots$ (1)
$\succ 2^{2}\cdot 3\succ 2^{2}\cdot 5\succ 2^{2}\cdot 7\succ\cdots\succ 2^{3}\succ 2^{2}\succ 2\succ 1.$
We write $m\succ n$ whenever $m$ is to the left of $n$ in Eq. (1). As a consequence of both theorems, a continuous one-dimensional map with period three has all periods. We note that such a map also has an uncountable set containing no (asymptotically) periodic points, the so-called scrambled set, and is chaotic in the sense of Li–Yorke [18].
Figure 1: Trained maps $f_{N}^{*}$ for activation functions $\phi=\text{erf}$ (top left), $\sin$ (top right), $\cos$ (bottom left), and ReLU (bottom right), with target period three ${\cal D}=\{-1,1,0\}$ and input intensity $\sigma=1.0$. The blue-colored areas indicate the maximum–minimum regions of $f_{N}^{*}$ for 100 different realizations. The red circles and the red dotted lines show ${\cal D}$ and the thermodynamic limit $f_{\infty}^{*}$, respectively. The darkness of blue corresponds to the number of nodes $N$, describing that $f_{N}^{*}$ degenerates into $f_{\infty}^{*}$ as $N$ increases.
Now, the following natural question arises: Can we obtain all the periods in the network through training only period three? If we successfully train period three in one-dimensional dynamics, then the straightforward answer is “yes.” However, this question is somewhat naive, since the above theorems do not reveal whether the obtained periods are stable. Instead, we should ask the following question: Which kind of stable orbits (attractors) can we obtain by learning period three? This paper is devoted to theoretically answer this question in terms of all aspects. Here, we consider training readout weights of a two-layer random feedforward neural network to learn period three [19, 20, 21, 22]: $f_{N}(x)\equiv\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{i}^{\text{out}}\phi\left(h_{i}(x)\right)$, $h_{i}(x)\equiv W_{i}^{\text{in}}x+b_{i}^{\text{in}}$, where $x\in\mathbb{R}$ is an input of the network, $W^{\text{in}}\in\mathbb{R}^{N\times 1}$ and $b^{\text{in}}\in\mathbb{R}^{N}$ are the input weights and biases, respectively, randomly drawn from a Gaussian distribution with zero mean and variance given by $\sigma^{2}$, $\phi\mathbin{:}\mathbb{R}\rightarrow\mathbb{R}$ is an activation function, and $W^{\text{out}}\in\mathbb{R}^{1\times N}$ is the output weights matrix optimized by a learning method described subsequently (see Sec. LABEL:S:sec:TL in the Supplementary Material [23] for the other network settings and detailed derivations, which includes Refs. [24, 25, 26]). We denote the trained network output by $f_{N}^{*}(x)$ and call $\sigma$ the input intensity. Hereafter, we consider the following specific activation functions: bounded and smooth $\phi=\text{erf},\sin,\cos$ and unbounded and non-smooth $\phi=\text{ReLU}$. Actually, this model can be regarded as a special case of an ESN in which the spectral radius of the adjacency matrix is zero. With the trained network $f_{N}^{*}$, we study the dynamical system $x_{n+1}=f_{N}^{*}(x_{n})$ created from a closed loop that connects its input and output. The attractors of $f_{N}^{*}$ should depend on the network structure, including the realizations of $W^{\text{in}}$ and $b^{\text{in}}$ and the choice of $\phi$, which makes the above question non-trivial.
Our findings are twofold. First, concerning general training data, we show that the trained map $f_{N}^{*}$ degenerates into its thermodynamic limit $f_{\infty}^{*}$ as $N$ increases (Fig. 1) based on the neural tangent kernel (NTK) theory [27, 28] under certain assumptions. Second, in learning period three, we reveal that in contrast to our expectation, almost all acquired periods are unstable; it has characteristic attractors corresponding to the choice of target period three, nonlinearity $\phi$, and input intensity $\sigma$. Interestingly, the changes in the target data points and $\sigma$, which can be regarded as controllable system parameters, cause the bifurcation of the characteristic attractors, enabling learning machines to embed a diverse range of dynamics in their bifurcation structure. We discuss the range of applicability of our proposed approach toward physical learning machines at the end of this paper.
We denote the training dataset and its number by ${\cal D}\subseteq\mathbb{R}\times\mathbb{R}$ and $|{\cal D}|$, respectively, and assume $N\geq|{\cal D}|$. We use ${\cal X}$ and ${\cal Y}$ vectors to denote the input and output data and define them as ${\cal X}\equiv[x_{1},\ldots,x_{|{\cal D}|}],{\cal Y}\equiv[y_{1},\ldots,y_{|{\cal D}|}]\in\mathbb{R}^{|{\cal D}|}$, where $(x_{i},y_{i})\in{\cal D}$. In learning period three ($a\mapsto b\mapsto c\mapsto a\mapsto\cdots$), the target input–output pairs are ${\cal D}=\left\{(a,b),(b,c),(c,a)\right\}$, ${\cal X}=[a,b,c]$, and ${\cal Y}=[b,c,a]$. For the sake of simplicity, we write ${\cal D}=\{a,b,c\}$ and assume $a<b$. Note that a period-3 orbit is of two types: $\{a,b,c\}$ and $\{b,a,c\}$. Generally, a period-$n$ orbit has $(n-1)!$ types. For a given ${\cal D}$, we optimize $W^{\text{out}}$ by minimizing a squared error via continuous-time gradient descent (GD) to easily take the limit $N\rightarrow\infty$ [27, 28, 29, 30].
We denote by $\theta_{t}\in\mathbb{R}^{N}$ and $f_{t}(x,\theta)\in\mathbb{R}$ the time-dependence of the vectorized output weights and $f_{N}(x)$ at time $t$, respectively. The dynamics of GD with the initialization $\theta_{0}=\bm{0}$ yield the following formulae of the optimal output weights $\left(W^{\text{out}}\right)^{*}$ and $f_{N}^{*}(x)$:
$\theta_{t}=\nabla_{\theta}f_{0}({\cal X})^{\top}\hat{\Theta}_{0}^{-1}\left(I-e^{-\eta\hat{\Theta}_{0}t}\right){\cal Y},$
$\left(W^{\text{out}}\right)^{*\top}=\lim_{t\rightarrow\infty}\theta_{t}=\nabla_{\theta}f_{0}({\cal X})^{\top}\hat{\Theta}_{0}^{-1}{\cal Y},$ (2)
$f_{t}(x,\theta)=\hat{\Theta}_{0}(x,{\cal X})\hat{\Theta}_{0}^{-1}\left(I-e^{-\eta\hat{\Theta}_{0}t}\right){\cal Y},$
$f_{N}^{*}(x)=\lim_{t\rightarrow\infty}f_{t}(x,\theta)=\hat{\Theta}_{0}(x,{\cal X})\hat{\Theta}_{0}^{-1}{\cal Y},$ (3)
where $\eta>0$ is a learning rate and $\nabla_{\theta}f_{0}({\cal X})\in\mathbb{R}^{|{\cal D}|\times N}$, $\hat{\Theta}_{0}(x,{\cal X})\in\mathbb{R}^{1\times|{\cal D}|}$, and $\hat{\Theta}_{0}\in\mathbb{R}^{|{\cal D}|\times|{\cal D}|}$ are the matrices; these matrices are given in the following manner:
$\left.\nabla_{\theta}f_{0}({\cal X})\right._{ij}=\frac{1}{\sqrt{N}}\phi\left(h_{j}(x_{i})\right),\,\left.\hat{\Theta}_{0}(x,{\cal X})\right._{i}\equiv\hat{\Theta}_{0}\left(x,x_{i}\right),$
$\left.\hat{\Theta}_{0}\right.\equiv\left.\hat{\Theta}_{0}({\cal X},{\cal X})\right.\equiv\nabla_{\theta}f_{0}({\cal X})\nabla_{\theta}f_{0}({\cal X})^{\top},$ (4)
$\hat{\Theta}_{0}(x,y)=\frac{1}{N}\sum_{i=1}^{N}\phi\left(h_{i}(x)\right)\phi\left(h_{i}(y)\right).$ (5)
We note that $\hat{\Theta}_{0}$ is symmetric and assumed to be positive definite. Eqs. (2) and (4) yield
$\left(W^{\text{out}}\right)^{*\top}=\nabla_{\theta}f_{0}({\cal X})^{\top}\left(\nabla_{\theta}f_{0}({\cal X})\nabla_{\theta}f_{0}({\cal X})^{\top}\right)^{-1}{\cal Y},$
implying that $\theta_{0}=\bm{0}$ will lead $W^{\text{out}}$ to the minimum norm solution (MNS) [29, 31, 30, 32], which is given by the pseudoinverse of $\nabla_{\theta}f_{0}({\cal X})$ (“extreme learning machine” [33, 20, 21]).
By the law of large numbers, $\hat{\Theta}_{0}(x,y)$ (Eq. (5)) converges in probability to $\Theta(x,y)$—that is the expectation over random variables $[\omega,\beta]\sim{\cal N}(\bm{0},\sigma^{2}I)$—in the limit $N\rightarrow\infty$ because the components of $W^{\text{in}}$ and $b^{\text{in}}$ follow an iid Gaussian distribution [27]:
[…]
$\Theta(x,y)$ is called the NTK [27]. Defining $\Theta(x,{\cal X})$ and $\Theta$ in the same manner as $\hat{\Theta}_{0}(x,y)$, we acquire $f^{*}_{\infty}(x)$, since $f_{N}^{*}(x)$ (Eq. (3)) is calculated from only the values of $\hat{\Theta}_{0}(x,y)$:
$f_{\infty}^{*}(x)=\Theta(x,{\cal X})\Theta^{-1}{\cal Y},$ (7)
where we again assume $\Theta$ to be positive definite. Moreover, the output of infinite-width ESN, which is described by recurrent kernels [34, 35, 36], reduces to the form of Eq. (7) in the limit where the regularization parameter goes to zero without considering the bias term of the network output.
With regard to learning period three, $f_{\infty}^{*}$ is given by
$f_{\infty}^{*}(x)=\begin{bmatrix}\Theta(x,a)\\ \Theta(x,b)\\ \Theta(x,c)\end{bmatrix}^{\top}\begin{bmatrix}\Theta(a,a)&\Theta(a,b)&\Theta(a,c)\\ \Theta(b,a)&\Theta(b,b)&\Theta(b,c)\\ \Theta(c,a)&\Theta(c,b)&\Theta(c,c)\end{bmatrix}^{-1}\begin{bmatrix}b\\ c\\ a\end{bmatrix}.$ (8)
To investigate the dynamics of $f_{\infty}^{*}$, we compute the trajectory $\left\{x_{n}\right\}_{n=0}^{T}$ of $T+1\gg 1$ steps with an initial state $x_{0}$ and calculate the Lyapunov exponent and period of attractors from trajectories. The Lyapunov exponent expresses the sensitivity of a dynamical system to initial conditions and is calculated using the following equation:
$\lambda_{T}=\frac{1}{T}\sum_{n=1}^{T}\ln{\left|\frac{d}{dx}f_{\infty}^{*}(x_{n})\right|}.$ (9)
We regard a trajectory of $f_{\infty}^{*}$ as chaotic when $\lambda_{T}>0$. Note that the derivative of $f_{\infty}^{*}$ is calculated by
$\frac{d}{dx}f_{\infty}^{*}(x)=\frac{\partial}{\partial x}\Theta(x,{\cal X})\Theta^{-1}{\cal Y}.$ (10)
We also calculate the period of attractors as the minimum integer $n\in[1,n_{\text{max}}]$ that satisfies the following inequality ($n_{\text{max}}=10\text{ or }20$):
$&\frac{\left|(f_{\infty}^{*})^{n}(x_{T})-x_{T}\right|}{\max\left\{\left|x_{T}\right|,\left|(f_{\infty}^{*})^{n}(x_{T})\right|\right\}}\leq 10^{-12},\\ \text{where}\quad&(f_{\infty}^{*})^{1}\equiv f_{\infty}^{*},\,(f_{\infty}^{*})^{k+1}\equiv f_{\infty}^{*}\circ(f_{\infty}^{*})^{k}.\end{aligned}\right.$ (11)
For $\phi=\text{erf},\sin,\cos,\text{ReLU}$, there exist the analytical solutions of $\Theta(x,y)$ (Sec. LABEL:S:subsec:NTK in the Supplementary Material [23], which includes Refs. [37, 38]):
$&\Theta^{\text{erf}}(x,y)\\ &=\frac{2}{\pi}\arcsin\frac{2\sigma^{2}(1+xy)}{\sqrt{\left[1+2\sigma^{2}(1+x^{2})\right]\left[1+2\sigma^{2}(1+y^{2})\right]}},\end{aligned}$ (12)
$\Theta^{\sin}(x,y)=\frac{1}{2}\left\{e^{-\frac{\sigma^{2}}{2}(x-y)^{2}}-e^{-\frac{\sigma^{2}}{2}(x+y)^{2}-2\sigma^{2}}\right\},$ (13)
$\Theta^{\cos}(x,y)=\frac{1}{2}\left\{e^{-\frac{\sigma^{2}}{2}(x-y)^{2}}+e^{-\frac{\sigma^{2}}{2}(x+y)^{2}-2\sigma^{2}}\right\},$ (14)
$\Theta^{\text{relu}}(x,y)=&\frac{\sigma^{2}}{2\pi}\sqrt{\left(1+x^{2}\right)\left(1+y^{2}\right)}\\ &\left\{\sqrt{1-\cos^{2}\psi}+(\pi-\psi)\cos\psi\right\},\end{aligned}$ (15)
where $\psi\equiv\arccos\frac{1+xy}{\sqrt{\left(1+x^{2}\right)\left(1+y^{2}\right)}}$. We note that the trajectories for a bounded activation function—such as $\phi=\text{erf},\sin,\cos$—are also bounded since $\Theta(x,y)$ (Eq. (6)) is the expectation of the product of $\phi$. Moreover, as $\phi=\text{erf},\sin,\cos$ and their corresponding NTKs are analytic, $f_{N}^{*}$ and $f_{\infty}^{*}$ have finitely many attractors, according to Theorem B in Ref. [39] (Sec. LABEL:S:sec:finiteness in the Supplementary Material [23], which includes Ref. [40]). In contrast, certain trajectories for $\phi=\text{ReLU}$, whose NTK is unbounded, will head toward infinity.
Figure 2: Change in dynamical system $f_{\infty}^{*}$ for $\phi=\text{erf}$ with respect to $c$, with $a=-1$, $b=1$, and $\sigma=1.0$. The trajectory $\{x_{n}\}_{n=0}^{T}$ of $T=10^{5}$ steps is computed with different $x_{0}$ and $c$ in given intervals, excluding $c=a,b$. (a) Bifurcation diagram of the characteristic attractors calculated with $-10\leq x_{0}\leq 10$. The dotted lines indicate ${\cal D}$; the diagonal line corresponds to a varying $c$. (b) Change in the basin of attraction where $f_{\infty}^{*}$ has multiple untrained attractors. (c) Change in the map $f_{\infty}^{*}$ in $-0.8\leq c\leq 0.8$. The circles indicate ${\cal D}$ and the red circle indicates a c-independent point $(a,b)=(-1,1)$. As $c$ approaches $a$ or $b$, the folding of $f_{\infty}^{*}$ around ${\cal D}$ becomes larger, making the characteristic attractors wider.
Figure 2 depicts how $f_{\infty}^{*}$ for $\phi=\text{erf}$ changes as $c$ varies. Note that the trained network always includes ${\cal D}$ since $f_{N}^{*}({\cal X})=f_{\infty}^{*}({\cal X})={\cal Y}$. In the case of ${\cal D}$ becoming the attractor ($-3\leq c<-2.2$ and $0<c\leq 0.22$), which corresponds to the successful attractor embedding by a learning machine, we observe that $f_{\infty}^{*}$ has no untrained attractors. However, varying $c$ changes the stability of all the periods—including ${\cal D}$—together, resulting in the emergence of untrained attractors: untrained period-3 ($-0.43\leq c<0$ and $2\leq c\leq 3$), chaotic ($c=-0.5,0.3$, etc.), and multiple attractors ($0.742\leq c\leq 0.764$, etc.). For $\phi=\sin,\cos,\text{ReLU}$, the bifurcation diagrams look very different from that in Fig. 2, but they have the same properties: only a handful of attractors appear at a time, which can be untrained ones (see Sec. LABEL:S:sec:dynamics in the Supplementary Material [23] for the bifurcation diagrams with the other network settings and the corresponding $\lambda_{T}$). We note that as the NTKs for $\phi=\text{erf},\sin,\cos,\text{ReLU}$ depend on the scaling and translation of inputs, varying $a$ or $b$, which we fixed, also yields a different bifurcation diagram. Nevertheless, the bifurcation structure has universal symmetry properties under the swapping of $a$ and $b$ (Sec. LABEL:S:sec:symmetry in the Supplementary Material [23]). In particular, there is a qualitative correspondence between the outside ($c<a,b<c$) and the inside ($a<c<b$) bifurcation structure, if $b\approx-a$ (Fig. 2(a)).
Figure 3: Two-dimensional slices of the characteristic bifurcation structures for $\phi=\text{erf}$ (top left), $\sin$ (top right), $\cos$ (bottom left), and ReLU (bottom right) with respect to $c$ and $\sigma$, with $a=-1$, $b=1$, $x_{0}=0$, and $T=10^{4}$. The black-hatched areas indicate the regions in which $\left|\frac{d}{dx}f_{\infty}^{*}(a)\frac{d}{dx}f_{\infty}^{*}(b)\frac{d}{dx}f_{\infty}^{*}(c)\right|<1$ holds. They correspond to the regions of ${\cal D}=\{a,b,c\}$ being locally stable, implying that for $\phi=\text{erf},\sin,\cos$, increasing $\sigma$ stabilizes ${\cal D}$. The blue-hatched area in the bottom-right panel indicates the region $(c,\sigma)$, where the trajectory heads toward infinity.
Figure 4: Two-dimensional slices of the characteristic bifurcation structure for $\phi=\text{erf}$ with respect to $c$ and $\sigma$, with $a=-1$, $b=1$, and $T=10^{4}$. (a) Number of kinds of period of attractors $N_{p}$ calculated with 100 equidistant initial states in $-10\leq x_{0}\leq 10$. $f_{\infty}^{*}$ has roughly $N_{p}$ attractors. (b) The zoom of (a). (c) The zoom of the top-left panel of Fig. 3. The black-hatched areas in (a),(b) indicate the regions of ${\cal D}=\{a,b,c\}$ being locally stable.
As we expect from the above discussion, we observe many types of characteristic attractors by varying $\sigma$ along with ${\cal D}$ (Figs. 3 and 4). Note that as $f_{\infty}^{*}$ may have multiple attractors (Fig. 4(a),(b)), certain trajectories reach attractors that are not depicted in Fig. 3. By utilizing these diagrams, we can control the untrained attractors of $f_{\infty}^{*}$ to obtain a large variety of attractors, which is extremely useful for generating intended dynamics. For example for $\phi=\text{erf}$, we obtain the period-13 attractor at $(c,\sigma)=(-0.665,1.32)$, period-19 attractor at $(c,\sigma)=(-0.655,1.172)$, and chaotic attractor at $(c,\sigma)=(-0.675,1.35)$ (“$+$” in Fig. 4(c)). For $\phi=\text{ReLU}$, the $\sigma$-dependence of NTK (Eq.(15)) is canceled out, resulting in the uniform bifurcation structure along the $\sigma$-direction (the bottom-right panel of Fig. 3). In contrast, for $\phi=\text{erf},\sin,\cos$, we can stabilize ${\cal D}$ by increasing $\sigma$ because the derivative of NTK at data point $y$ approaches zero for large $\sigma$ (the black-hatched areas in Fig. 3):
$\frac{\partial}{\partial x}\Theta^{\text{erf}}(y,y)=\frac{4\sigma^{2}y}{\pi\left[1+2\sigma^{2}(1+y^{2})\right]\sqrt{1+4\sigma^{2}(1+y^{2})}},$
$\frac{\partial}{\partial x}\Theta^{\sin}(y,y)=-\frac{\partial}{\partial x}\Theta^{\cos}(y,y)=\sigma^{2}ye^{-2\sigma^{2}(1+y^{2})}.$
However, $f_{\infty}^{*,\sin}$ and $f_{\infty}^{*,\cos}$ exponentially decrease to zero except at ${\cal D}$ as $\sigma$ grows; perturbations of $x_{0}$ from ${\cal D}$ lead to their different attractors for large $\sigma$; $\Theta^{\sin}(x,y)$ and $\Theta^{\cos}(x,y)$ become the Kronecker-delta-like functions in the infinite input-intensity limit $\sigma\rightarrow\infty$:
$\lim_{\sigma\rightarrow\infty}\Theta^{\sin}(x,y)=\lim_{\sigma\rightarrow\infty}\Theta^{\cos}(x,y)=\frac{1}{2}\bm{1}_{y}(x),$ (16)
where $\bm{1}_{y}(x)\equiv 1$ for $x=y$ and 00 otherwise. Interestingly, for $\phi=\text{erf}$, $\Theta^{\text{erf}}(x,y)$ and, therefore, $f_{\infty}^{*,\text{erf}}$ becomes piecewise-monotonic and piecewise-smooth, with ${\cal D}$ being its singular points in the limit $\sigma\rightarrow\infty$:
$\lim_{\sigma\rightarrow\infty}\Theta^{\text{erf}}(x,y)=\frac{2}{\pi}\arcsin\frac{1+xy}{\sqrt{(1+x^{2})(1+y^{2})}},$ (17)
$\frac{\partial}{\partial x}\lim_{\sigma\rightarrow\infty}\Theta^{\text{erf}}(x,y)=\frac{-2}{\pi(1+x^{2})}\cdot\frac{x-y}{|x-y|},$ (18)
enabling the candidates of robust chaos [41, 42] to appear. In addition, in this limit, NTK for sigmoidal activation (17) is equivalent to that for binary activation ($\phi=\text{sgn}$) [36]. Although $f_{N}^{*}$ for $\phi=\text{sgn}$ is beyond the scope of theorems 1 and 2, it asymptotically approaches a continous map $f_{\infty}^{*}$, exhibiting multiple stable periodic orbits for large $N$. Finite-size effects also emerge for $\phi=\sin,\cos$, which are caused by the wavy deviations of $f_{N}^{*}$ from $f_{\infty}^{*}$ (Sec. LABEL:S:sec:dynamics in the Supplementary Material [23]).
Finally, we discuss the peculiarity of learning period three with MNS and the scalability of our thermodynamic limit analysis in dynamical systems learning. Since $f_{\infty}^{*}(x)$ is described by the multiplication of $\Theta(x,{\cal X})$ and $\Theta^{-1}$, $|{\cal D}|$ crucially affects the complexity of the theoretical analysis of $f_{\infty}^{*}$. Accordingly, the longer the target orbit becomes—such as in the case of learning chaos and learning continuous-time dynamical systems from time series—the fewer the benefits of our proposed approach. However, similar to learning period three, if we adopt the strategy of exploiting the intrinsic mathematical structure of a learning machine with a small amount of training data, infinite-width networks provide a novel perspective to dynamical systems learning. We also investigated how the learned periods depend on the target period number $n$ and found that regardless of the choice of $\phi$, the number of unstable periods tends to increase dramatically after $n=3$ (Sec. LABEL:S:sec:LPn in the Supplementary Material [23]). This result suggests that a rich characteristic bifurcation structure is obtained by a combination of bulk unstable periods ensured by learning period three and a change in their stability through the system’s nonlinearity, which is our answer to the very first question.
As an engineering application, our results suggest that various types of dynamics, including higher-order periodic orbits and chaos, can be obtained by teaching only three target data points to the network with appropriate settings. More importantly, the network may be a physical system or neuromorphic device, such as the physical reservoir computer [43]; embedding desirable dynamics into physical systems would be highly beneficial, for example, in a high-radiation environment in which a computer-simulated attractor may break down [44]. We leave the thermodynamic limit analysis of a physical neural network for future work.
Acknowledgements.
We are grateful to Allen Hart for the fruitful discussions on attractor embedding in RC and to the RC seminar members for the stimulating discussions.