### v2

# Thermodynamic Limit in Learning Period Three

Abstract
A continuous one-dimensional map with period three includes all periods. This raises the following question: Can we obtain any types of periodic orbits solely by learning three data points? In this letter, we report the answer to be yes. Considering a random neural network in its thermodynamic limit, we show that under certain conditions, learning period three can embed attractors with all periods into the network as a bifurcation after learning. The associated universality is explained by a topological conjugacy between the trained network and the classical logistic map.
With the advent of reservoir computing (RC) [1], which exploits the dynamics of high-dimensional dynamical systems for learning, many interesting properties of random networks have been discovered from a dynamical systems perspective. For example, with a fixed-weight random recurrent neural network, the so-called echo-state network (ESN) [2]—as a reservoir—we can create an autonomous system that emulates target chaotic systems by simply fitting its readout layer [3, 4]. Such an attractor-embedding ability is linked to the existence of the generalized synchronization between an input dynamical system and a reservoir [4, 5, 6, 7]. In addition, recent studies have revealed that RC can simultaneously embed multiple attractors [8] and may have untrained attractors that are not part of the training data [8, 9, 10, 11, 12, 13, 14, 15]. Surprisingly, a successful reservoir computer only needs a few pairs of bifurcation parameter values and corresponding trajectories to reconstruct the entire bifurcation structure of a target system [10, 11, 12, 13, 14]. These properties are valuable, for example, in the context of robot locomotion control using dynamical system attractors that can significantly reduce the training data [16, 17, 18]. Thus, the powerful generalization and multifunctionality aspects of RC are related to the dynamical systems properties of a learning machine.
In one-dimensional discrete dynamical systems, there are two significant theorems on periodic orbits [19, 20, 21]:
Theorem 1
(Sharkovsky) If a continuous map $f:I\rightarrow I$ has a periodic point of period $m$ and $m\succ n$, then $f$ also has a periodic point of period $n$.
Theorem 2
(Li–Yorke) If a continuous map $f:I\rightarrow I$ has a point $a\in I$ for which the points $b=f(a)$, $c=f(f(a))$, and $d=f(f(f(a)))$ satisfy
$d\leq a<b<c\quad({\rm or}\quad d\geq a>b>c),$
then $f$ has a periodic point of period $k$ for every $k\in\mathbb{N}$.
Note that the interval $I$ does not need to be closed or bounded. Here, the ordering of positive integers $\succ$ in Theorem 1 is called the Sharkovsky ordering and is given below:
33 33 $\succ 5\succ 7\succ 9\succ\cdots\succ 2\cdot 3\succ 2\cdot 5\succ 2\cdot 7\succ\cdots$ (1)
$\succ 2^{2}\cdot 3\succ 2^{2}\cdot 5\succ 2^{2}\cdot 7\succ\cdots\succ 2^{3}\succ 2^{2}\succ 2\succ 1.$
We write $m\succ n$ whenever $m$ is to the left of $n$ in Eq. (1). As a consequence of both theorems, a continuous one-dimensional map with period three has all periods.
Now, the following natural question arises: Can we obtain all the periods in the network through training only period three? If we successfully train period three in one-dimensional dynamics, then the straightforward answer is “yes.” However, this question is somewhat naive, since the above theorems do not reveal whether the obtained periods are stable. Instead, we should ask the following question: Which kind of stable orbits (attractors) can we obtain by learning period three (LP3)? This paper is devoted to theoretically answer this question in terms of all aspects.
To clarify what LP3 implies, we need to create an one-dimensional recurrent neural network with target period three. LP3 with ESN, as in the standard RC scheme, will violate the assumption of the above two theorems due to the high-dimensionality of the resulting dynamics. Thus, here, we consider training the readout weights of a two-layer random feedforward neural network to learn period three [22, 23, 24, 25]: $f_{N}(x)\equiv\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{i}^{\text{out}}\phi\left(h_{i}(x)\right)$, $h_{i}(x)\equiv W_{i}^{\text{in}}x+b_{i}^{\text{in}}$, where $x\in\mathbb{R}$ is an input of the network; $W^{\text{in}}\in\mathbb{R}^{N\times 1}$ and $b^{\text{in}}\in\mathbb{R}^{N}$ are the input weights and biases, respectively, randomly drawn from a Gaussian distribution with zero mean and variance given by $\sigma^{2}$; $\phi\mathbin{:}\mathbb{R}\rightarrow\mathbb{R}$ is an activation function; and $W^{\text{out}}\in\mathbb{R}^{1\times N}$ is the output weights matrix optimized by a learning method described subsequently (see Sec. LABEL:S:subsec:derivation and Sec. LABEL:S:subsec:uniform in the Supplementary Material (SM) [26] for networks with output bias and uniform distribution). Actually, this model can be regarded as a special case of an ESN in the limit where the spectral radius of the adjacency matrix goes to zero. In addition, we denote the trained network output by $f_{N}^{*}(x)$ and call $\sigma$ the scale of weights. With the trained network $f_{N}^{*}$, we study the dynamical system $x_{n+1}=f_{N}^{*}(x_{n})$ created from a closed loop that connects its input and output. The attractors of $f_{N}^{*}$ should depend on the network structure, including the realizations of $W^{\text{in}}$ and $b^{\text{in}}$ and the choice of $\phi$, which makes the above question non-trivial. Hereafter, we consider the following specific activation functions: bounded and smooth $\phi=\text{erf},\sin,\cos$, and unbounded and non-smooth $\phi=\text{ReLU}$.
Figure 1: Trained maps $f_{N}^{*}$ for activation functions $\phi=\text{erf}$ (top left), $\sin$ (top right), $\cos$ (bottom left), and ReLU (bottom right), with target period three ${\cal D}=\{-1,1,0\}$ and scale of weights $\sigma=1.0$. The blue-colored areas indicate the maximum–minimum regions of $f_{N}^{*}$ for 100 different realizations. The red circles and the red dotted lines show ${\cal D}$ and the thermodynamic limit $f_{\infty}^{*}$, respectively. The darkness of blue corresponds to the number of nodes $N$, describing that $f_{N}^{*}$ degenerates into $f_{\infty}^{*}$ as $N$ increases.
Our findings are threefold. First, concerning general training data, we show that the trained map $f_{N}^{*}$ degenerates into its thermodynamic limit $f_{\infty}^{*}$ as $N$ increases (Fig. 1) under certain assumptions. Second, in LP3, we reveal that almost all acquired periods are unstable; it has characteristic attractors corresponding to the choice of target period three, nonlinearity $\phi$, and scale of weights $\sigma$. Third, we prove that neural networks with certain activations can externalize all the latently acquired unstable periods through LP3 as attractors after learning, as a universal bifurcation with regard to the feedback strength. We discuss the range of applicability of our proposed approach to physical learning machines at the end of this paper.
We denote the training dataset and its number by ${\cal D}\subseteq\mathbb{R}\times\mathbb{R}$ and $|{\cal D}|$, respectively, and assume $N\geq|{\cal D}|$. We use ${\cal X}$ and ${\cal Y}$ vectors to denote the input and output data and define them as ${\cal X}\equiv[x_{1},\ldots,x_{|{\cal D}|}],{\cal Y}\equiv[y_{1},\ldots,y_{|{\cal D}|}]\in\mathbb{R}^{|{\cal D}|}$, where $(x_{i},y_{i})\in{\cal D}$. In LP3 ($a\mapsto b\mapsto c\mapsto a\mapsto\cdots$), the target input–output pairs are ${\cal D}=\left\{(a,b),(b,c),(c,a)\right\}$, ${\cal X}=[a,b,c]$, and ${\cal Y}=[b,c,a]$. For the sake of simplicity, we write ${\cal D}=\{a,b,c\}$ and assume $a<b$. Note that a period-3 orbit is of two types: $\{a,b,c\}$ and $\{b,a,c\}$. Generally, a period-$n$ orbit has $(n-1)!$ types. For a given ${\cal D}$, we optimize $W^{\text{out}}$ by least square regression with a minimum norm solution (“extreme learning machine” [27, 23, 24]) or “ridgeless” regression [28, 29, 30, 31]:
$f_{N}^{*}(x)=\lim_{\lambda\searrow 0}\hat{\Theta}(x,{\cal X})\left(\hat{\Theta}+\lambda I_{|{\cal D}|}\right)^{-1}{\cal Y},$ (2)
where $\hat{\Theta}(x,{\cal X})\in\mathbb{R}^{1\times|{\cal D}|}$ and $\hat{\Theta}\equiv\hat{\Theta}({\cal X},{\cal X})\in\mathbb{R}^{|{\cal D}|\times|{\cal D}|}$ are the matrices given in the following manner:
$\left.\hat{\Theta}(x,{\cal X})\right._{i}\equiv\hat{\Theta}\left(x,x_{i}\right),\left.\hat{\Theta}\right._{ij}\equiv\hat{\Theta}\left(x_{i},x_{j}\right),$ (3)
$\hat{\Theta}(x,y)\equiv{\cal R}(x)^{\top}{\cal R}(y)=\frac{1}{N}\sum_{i=1}^{N}\phi\left(h_{i}(x)\right)\phi\left(h_{i}(y)\right),$ (4)
${\cal R}(x)_{i}\equiv\frac{1}{\sqrt{N}}\phi\left(h_{i}(x)\right),\quad{\cal R}(x)\in\mathbb{R}^{N}.$ (5)
We note that Eq. (2) is equivalent to fit $W^{\text{out}}$ with the pseudoinverse of the matrix of hidden states ${\cal R}({\cal X})\equiv[{\cal R}(x_{1})\cdots{\cal R}(x_{|{\cal D}|})]\in\mathbb{R}^{N\times|{\cal D}|}$ [27]: $\left(W^{\text{out}}\right)^{*}={\cal Y}^{\top}{\cal R}({\cal X})^{+}$. If the matrix $\hat{\Theta}$ has full rank, Eq. (2) has the following closed-form expression:
$f_{N}^{*}(x)=\hat{\Theta}(x,{\cal X})\hat{\Theta}^{-1}{\cal Y}.$ (6)
By the law of large numbers, $\hat{\Theta}(x,y)$ (Eq. (4)) converges in probability to $\Theta(x,y)$—that is the expectation over random variables $[\omega,\beta]\sim{\cal N}(\bm{0},\sigma^{2}I)$—in the limit $N\rightarrow\infty$ because the components of $W^{\text{in}}$ and $b^{\text{in}}$ follow an iid Gaussian distribution [32]:
[…]
In our model, $\Theta(x,y)$ coincides with the neural tangent kernel (NTK) and the neural network gaussian process (NNGP) kernel [32, 33]. Defining $\Theta(x,{\cal X})$ and $\Theta$ in the same manner as $\hat{\Theta}(x,y)$, we acquire $f^{*}_{\infty}(x)$, since $f_{N}^{*}(x)$ (Eq. (6)) is calculated from only the values of $\hat{\Theta}(x,y)$:
$f_{\infty}^{*}(x)=\Theta(x,{\cal X})\Theta^{-1}{\cal Y},$ (8)
where we again assume $\Theta$ to have full rank. This assumption is valid if $\phi$ is a non-polynomial continuous function and ${\cal X}$ consists of $|{\cal D}|$ distinct points, according to Theorem 3 in Ref. [34] (Sec. LABEL:S:subsec:gram in the SM [26], which includes Ref. [35]). In addition, the output of the infinite-width ESN also reduces to a form similar to that of Eq. (8) with time-varying recurrent kernels [36, 37, 38].
With regard to LP3, $f_{\infty}^{*}$ is given by
$f_{\infty}^{*}(x)=\begin{bmatrix}\Theta(x,a)\\ \Theta(x,b)\\ \Theta(x,c)\end{bmatrix}^{\top}\begin{bmatrix}\Theta(a,a)&\Theta(a,b)&\Theta(a,c)\\ \Theta(b,a)&\Theta(b,b)&\Theta(b,c)\\ \Theta(c,a)&\Theta(c,b)&\Theta(c,c)\end{bmatrix}^{-1}\begin{bmatrix}b\\ c\\ a\end{bmatrix}.$ (9)
To investigate the dynamics of $f_{\infty}^{*}$, we compute the trajectory $\left\{x_{n}\right\}_{n=0}^{T}$ of $T+1\gg 1$ steps with an initial state $x_{0}$ and calculate the Lyapunov exponent and period of attractors from the trajectories. The Lyapunov exponent expresses the sensitivity of a dynamical system to initial conditions and is calculated using the following equation:
$\lambda_{T}=\frac{1}{T}\sum_{n=1}^{T}\ln{\left|\frac{df_{\infty}^{*}}{dx}(x_{n})\right|}.$ (10)
We regard a trajectory of $f_{\infty}^{*}$ as chaotic when $\lambda_{T}>0$. Note that the derivative of $f_{\infty}^{*}$ is calculated by
$\frac{df_{\infty}^{*}}{dx}(x)=\frac{\partial\Theta}{\partial x}(x,{\cal X})\Theta^{-1}{\cal Y}.$ (11)
We also calculate the period of attractors as the minimum integer $n\in[1,n_{\text{max}}]$ that satisfies the following inequality ($n_{\text{max}}=10\text{ or }20$ and $\varepsilon=10^{-12}$ except as otherwise noted):
$&\frac{\left|(f_{\infty}^{*})^{n}(x_{T})-x_{T}\right|}{\max\left\{\left|x_{T}\right|,\left|(f_{\infty}^{*})^{n}(x_{T})\right|\right\}}\leq\varepsilon,\\ \text{where}\quad&(f_{\infty}^{*})^{1}\equiv f_{\infty}^{*},\,(f_{\infty}^{*})^{k+1}\equiv f_{\infty}^{*}\circ(f_{\infty}^{*})^{k}.\end{aligned}\right.$ (12)
For $\phi=\text{erf},\sin,\cos,\text{ReLU}$, there exist the analytic solutions of $\Theta(x,y)$. Following Williams [39], the NTK for $\phi=\text{erf}$ is expressed as follows (see Sec. LABEL:S:subsec:NTK in the SM [26] regarding the NTKs for $\phi=\sin,\cos,\text{ReLU}$, which includes Ref. [40]):
$\Theta^{\text{erf}}(x,y)$ (13)
$=\frac{2}{\pi}\arcsin\frac{2\sigma^{2}(1+xy)}{\sqrt{\left[1+2\sigma^{2}(1+x^{2})\right]\left[1+2\sigma^{2}(1+y^{2})\right]}}.$
We note that the trajectories for a bounded activation function—such as $\phi=\text{erf},\sin,\cos$—are also bounded since $\Theta(x,y)$ (Eq. (7)) is the expectation of the product of $\phi$. Moreover, as $\phi=\text{erf},\sin,\cos$ and their corresponding NTKs are analytic, $f_{N}^{*}$ and $f_{\infty}^{*}$ have finitely many attractors, according to Theorem B in Ref. [41] (Sec. LABEL:S:sec:finiteness in the SM [26], which includes Ref. [42]). In contrast, certain trajectories for $\phi=\text{ReLU}$, whose NTK is unbounded, will head toward infinity.
Figure 2: Change in dynamical system $f_{\infty}^{*}$ for $\phi=\text{erf}$ with respect to $c$, with $a=-1$, $b=1$, and $\sigma=1.0$. The trajectory $\{x_{n}\}_{n=0}^{T}$ of $T=10^{5}$ steps is computed with different $x_{0}$ and $c$ in given intervals, excluding $c=a,b$. (a) Changes in the acquired periods $\{x_{p}\}$ of period $p$, calculated by solving $(f_{\infty}^{*})^{p}(x_{p})=x_{p}$ (see Sec. LABEL:S:sec:LPn in the SM [26] for details). (b) Bifurcation diagram of the characteristic attractors calculated with $-10\leq x_{0}\leq 10$. The dotted lines indicate ${\cal D}$; the diagonal lines correspond to varying $c$. (c) Change in the basin of attraction where $f_{\infty}^{*}$ has multiple untrained attractors. (d) Change in the map $f_{\infty}^{*}$ in $-0.8\leq c\leq 0.8$. The circles indicate ${\cal D}$ and the red circle indicates a c-independent point $(a,b)=(-1,1)$. As $c$ approaches $a$ or $b$, the folding of $f_{\infty}^{*}$ around ${\cal D}$ becomes larger, making the characteristic attractors wider.
Figure 2 depicts how $f_{\infty}^{*}$ for $\phi=\text{erf}$ changes as $c$ varies. Note that as long as $\hat{\Theta}$ and $\Theta$ have full rank, trained networks completely learn the target orbit ${\cal D}$ since $f_{N}^{*}({\cal X})=f_{\infty}^{*}({\cal X})={\cal Y}$. In the case of ${\cal D}$ becoming the attractor ($-3\leq c<-2.2$ and $0<c\leq 0.22$), which corresponds to the successful attractor embedding by a learning machine, we observe that $f_{\infty}^{*}$ has no untrained attractors. However, varying $c$ causes the bifurcation of the embeddable attractors of $f_{\infty}^{*}$, resulting in the emergence of untrained attractors: untrained period-3 ($-0.43\leq c<0$ and $2\leq c\leq 3$), chaotic ($c=-0.5,0.3$, etc.), and multiple attractors ($0.742\leq c\leq 0.764$, etc.). Consequently, only a handful of attractors appear at a time, and almost all periods latently exist as unstable periods (Fig. 2(a),(b)). Such bifurcation of the characteristic attractors strongly depends on the network settings (see Sec. LABEL:S:sec:dynamics in the SM [26] for the bifurcation diagrams with respect to $c$ and $\sigma$, along with the corresponding $\lambda_{T}$, which includes Refs. [43, 44]). We note that as the NTKs for $\phi=\text{erf},\sin,\cos,\text{ReLU}$ depend on the scaling and translation of inputs, varying $a$ or $b$, which we fixed, also yields a different bifurcation. Nevertheless, the bifurcation structure has universal symmetry properties under the swapping of $a$ and $b$ (Sec. LABEL:S:sec:symmetry in the SM [26]). In particular, if $\Theta(x,y)=\Theta(-x,-y)$, $\left.f_{\infty}^{*}\right|_{{\cal D}=\{a,-a,c\}}$ is topologically conjugate to $\left.f_{\infty}^{*}\right|_{{\cal D}=\{-a,a,-c\}}$, and there is a qualitative correspondence between the outside ($c<a,b<c$) and the inside ($a<c<b$) bifurcation structure, if $b\approx-a$ (Fig. 2(a),(b)).
Figure 3: Change in dynamical system $\sigma_{\text{fb}}f_{\infty}^{*}$ for $\phi=\text{erf}$ with respect to feedback strength $\sigma_{\text{fb}}$, with ${\cal D}=\{-1,1,0\}$ and $\sigma=1$. (a) Bifurcation diagram calculated with $-10\leq x_{0}\leq 10$ and $T=10^{5}$. (b) Stability changes in the acquired periods $\{x_{p}\}$ of period $p$, calculated by $\lambda_{p}\equiv\ln\left|\frac{d}{dx}(\sigma_{\text{fb}}f_{\infty}^{*})^{p}(x_{p})\right|$ using the solutions of $(\sigma_{\text{fb}}f_{\infty}^{*})^{p}(x_{p})=x_{p}$. The vertical dotted lines correspond to the network state in LP3 ($\sigma_{\text{fb}}=1$). The horizontal dotted line indicates the boundary of the stability $\lambda_{p}=0$; $\lambda_{p}<0$ and $\lambda_{p}>0$ mean (locally) stable and unstable, respectively. In this setting, decreasing $\sigma_{\text{fb}}$ toward 0 externalizes the latently existing unstable periods as period-doubling cascades, which is illustrated as the continuous lines of $\{\lambda_{p}\}$ from $\lambda_{p}>0$ (at $\sigma_{\text{fb}}=1$) to $\lambda_{p}<0$ (at some $\sigma_{\text{fb}}\in(0,1)$).
So far, we have only focused on the characteristic attractors of the neural networks in LP3, which are just small parts of latently existing infinitly many periodic orbits. To illuminate the meaning of the latently existing unstable periods, we further extend our learning procedure to varying the readout weights of the trained neural networks [25]. Figure 3 depicts how its attractor changes as the scale of the readout weights—the feedback strength $\sigma_{\text{fb}}$—varies. We set the network state of LP3 to $\sigma_{\text{fb}}=1$; this dynamical system is expressed as $\sigma_{\text{fb}}f_{\infty}^{*}$. Since the network is prohibited from having any periodic orbits except a fixed point $x=0$ at $\sigma_{\text{fb}}=0$, all the acquired periods $\{x_{p}\}$ must disappear at some $\sigma_{\text{fb}}\in(0,1)$ with $\frac{d}{dx}(\sigma_{\text{fb}}f_{\infty}^{*})^{p}(x_{p})=1$ (tangent bifurcation) due to the differentiability of $f_{\infty}^{*}$; some of them may become attractors before dying out. The way of the externalization of latently unstable periods depends on the choice of ${\cal D}$ and the network structure (see Fig. LABEL:S:fig:externalize_inf in the SM [26]).
Now, let us assume that ${\cal D}=\{a,-a,0\}$ (symmetric LP3) and $\phi=\text{erf},\sin,\cos$. Surprisingly, this symmetry of ${\cal D}$ and differentiability of $\phi$ will induce further universal bifurcation that externalizes all the latently unstable periods into stable ones after learning, corresponding to the Sharkovsky ordering (Eq. (1)) (see Figs. LABEL:S:fig:externalize_universal_infLABEL:S:fig:externalize_universal in the SM [26] for $\phi=\text{ReLU}$, which induces slightly different externalization). Thanks to the symmetry in the distribution of $W^{\text{in}}$ at $W^{\text{in}}=0$, $\Theta(x,x),\Theta(x,-x)$, and $\Theta(x,0)$ are even functions of $x$ since $\Theta(x,y)=\Theta(-x,-y)$. Applying $b=-a$ and $c=0$ to Eq. (9) and exploiting these symmetries, the derivative of $f_{\infty}^{*}$ at $x=0$ is given as follows:
$\left.\frac{df_{\infty}^{*}}{dx}(0)\right|_{{\cal D}=\{a,-a,0\}}=-\frac{a\frac{\partial\Theta}{\partial x}(0,a)}{\Theta(a,a)-\Theta(a,-a)}$ (14)
$=-\frac{\frac{\partial^{2}\Theta}{\partial x\partial y}(0,0)+{\cal O}(a)}{2\frac{\partial^{2}\Theta}{\partial x\partial y}(0,0)+{\cal O}(a)}\rightarrow-\frac{1}{2}\quad(a\rightarrow 0).$
Note that $\frac{\partial^{2}\Theta}{\partial x\partial y}(0,0)=\mathbb{E}_{\omega,\beta}\left[\omega^{2}\phi^{\prime}(\beta)^{2}\right]>0$. Next, for $a\ll 1$, the local shape of $f_{\infty}^{*}(x)$ around ${\cal D}=\{a,-a,0\}$ is well described as a polynomial function $g(x)$. Imposing $g({\cal X})={\cal Y}$ and $\frac{dg}{dx}(0)=-\frac{1}{2}$, its third-order approximation is uniquely determined as
$g(x)=a\hat{g}\left(\frac{x}{a}\right),\quad\hat{g}(x)\equiv 1-\frac{x}{2}-\frac{3x^{2}}{2}+0\cdot x^{3},$ (15)
which is topologically conjugate to $\hat{g}$. Furthermore, under $\mu>2$ and $\sigma_{\text{fb}}>0$, $x_{n+1}=\sigma_{\text{fb}}\hat{g}(x_{n})$ is topologically conjugate to the logistic map $y_{n+1}=\mu y_{n}(1-y_{n})$ in the following manner [45]:
$x_{n}=-\frac{2}{3}\left[\frac{\mu}{\sigma_{\text{fb}}(\mu)}\left(-y_{n}+\frac{1}{2}\right)+\frac{1}{4}\right],$ (16)
$\sigma_{\text{fb}}(\mu)=\frac{4}{5}\sqrt{\frac{\mu}{2}\left(\frac{\mu}{2}-1\right)+\frac{1}{100}}-\frac{2}{25}.$ (17)
With $0<\sigma_{\text{fb}}(\mu_{1})\leq\sigma_{\text{fb}}(\mu_{2})$ for $2<\mu_{1}\leq\mu_{2}$ and the Sharkovsky ordering in $\mu$ [46, 47], we obtain the Sharkovsky ordering in externalization:
$\sigma_{\text{fb}}[3]$ $\geq\sigma_{\text{fb}}[5]\geq\cdots\geq\sigma_{\text{fb}}[2\cdot 3]\geq\sigma_{\text{fb}}[2\cdot 5]\geq\cdots$ (18)
$\geq\sigma_{\text{fb}}[2^{2}\cdot 3]\geq\cdots\geq\sigma_{\text{fb}}[2^{2}]\geq\sigma_{\text{fb}}[2]\geq\sigma_{\text{fb}}[1],$
where we denote by $\sigma_{\text{fb}}[n]$ the least value of $\sigma_{\text{fb}}$ for which $\sigma_{\text{fb}}\hat{g}$ possesses period $n$ as its attractor and $\sigma_{\text{fb}}[1]\equiv 0$. Here, $10=\sigma_{\text{fb}}[1]\leq\cdots\leq\sigma_{\text{fb}}[3]=\sigma_{\text{fb}}(1+\sqrt{8})<$ 1. Consequently, the universality (Eqs. (15) and (18)) retains no information about network settings $N,\phi,\sigma$ or the dataset $a$, implying the strong convergence of the externalization in symmetric LP3 (see Sec. LABEL:S:sec:externalize in the SM [26] for a detailed discussion, which includes Ref. [48]). Interestingly, with a small $a$, we observed the coexistence of the controlled attractors—which come from the universality $\hat{g}$—and the characteristic attractors—which come from the details of the network structure—in the externalization $\sigma_{\text{fb}}f_{\infty}^{*}$ (see Fig. LABEL:S:fig:externalize_universal_inf in the SM [26]).
Although LP3 is not a necessary condition for achieving period three (learning period two or even random neural networks [49] may have period three, see Fig. LABEL:S:fig:period_n_distribution and Fig. LABEL:S:fig:randomNN in the SM [26]), it provides a sufficient condition for embedding attractors with all periods as a universal bifurcation along with externally controllable parameters ${\cal D}$ and $\sigma_{\text{fb}}$—which is our answer to the very first question. LP3 also provides new yet important perspectives in the learning of dynamics. First, even if neural networks completely learn a target orbit ${\cal D}$, they may fail to replicate ${\cal D}$, since its stability depends on the local structure of the trained map $f_{\infty}^{*}$ around ${\cal D}$. Second, LP3 is not a goal, but a groundwork or a primer for updating the connectivity of random networks to generate all types of periodic orbits, including chaos, after learning. Generating network dynamics with a minimal dataset is not just efficient but also compatible with the theoretical analysis, since $f_{\infty}^{*}(x)$ is described by the multiplication of $\Theta(x,{\cal X})$ and $\Theta^{-1}$; learning dynamics from time series lead to a large $|{\cal D}|$ and a nearly singular $\Theta$, making the analysis of $f_{\infty}^{*}$ ineffective.
As an engineering application, we may replace the feedforward network part with a physical system or neuromorphic device [50, 51, 52]; embedding desirable dynamics into physical systems would be highly beneficial, for example, in a high-radiation environment in which a computer-simulated attractor may break down [53]. We leave the further analysis on the control of embeddable attractors in physical systems for future work.
Acknowledgements.
We are grateful to Allen Hart for the fruitful discussions on attractor embedding in RC and to the RC seminar members for the stimulating discussions.